My mind has gone blank....any help is greatly appreciated:
A boat has a speed of 5 mph in still water. The boat can travel 21 miles with the current in the same time in which it can travel 9 miles against the current. Find the rate of the current.

Answer by edjones(8007)

d=r*t
t=d/r
Let x be the rate of the current.
21/(5+x)=9/(5-x)
21(5-x)=9(5+x) Multiply each side by the LCM to eliminate fractions.
105-21x=45+9x
60=30x
x=2 mph
.
Check:
21/7=3 hrs
9/3=3 hrs
.
Ed